Status Report

NASA Space Station On-Orbit Status 30 July 2005

By SpaceRef Editor
July 30, 2005
Filed under , , ,
NASA Space Station On-Orbit Status 30 July 2005

SpaceRef note: This NASA Headquarters internal status report, as presented here, contains additional, original material produced by (copyright © 2005) to enhance access to related status reports and NASA activities.

All ISS systems continue to function nominally, except those noted previously or below.  Flight Day 5 (FD5) of the STS-114/LF-1 mission.  Today Sergei Krikalev has accumulated 2 years of time spent in space. Congratulations, Sergei!

EVA-1 accomplished all objectives.

After crew wakeup at the current 12:10am EDT this morning, CDR Krikalev and FE/SO Phillips spent most of the day supporting EVA-1, the first spacewalk (of three planned) of Shuttle Mission Specialists Steve Robinson and Soichi Noguchi.

Processing Status
Daily Mission
Return to Flight
Weekly Status
Weekly Science
Daily On-Orbit Status
Daily Crew Timeline
Soyuz | Progress
ISS News | ATV

The EVA began at 5:46am, about an hour later than planned due preparation and suiting-up delays.  The spacewalkers, egressing from the Shuttle airlock and keeping the ISS Airlock crewlock (A/L C/L) as backup point of ingress, successfully accomplished all tasks.   [These included TPS DTO (thermal protection system development test objective), ESP2 (External Stowage Platform 2) cable routing for platform heater power, ESPAD (ESP Attachment Device) removal from the Shuttle cargo bay and installation on the ISS A/L, removal and replacement (R&R) of GPS antenna AA2 on the S0 truss, Z1 patch panel reconfiguration for CMG-2 (control moment gyroscope 2) activation and operation, and CMG-3 R&R get-aheads.  The ESP2 platform itself and its supply of spare components will be relocated from the Shuttle to the ISS during EVA-3.  Initial power-up and spin-up of CMG-2 was successful.  Full spin-up and addition of the CMG to the vehicle attitude management system will be completed tomorrow, during FD6.  The GPS AA2 antenna was confirmed as fully functional.]   

The TPS DTO activities in the Shuttle payload bay began at ~6:45am and extended thru 8:16am EDT.   [Preliminary test results:  The NOAX (Non-Oxide Adhesive eXperimental) demo on the RCC porosity sample showed some swelling after its application to cracks.  When smoothed out with the spatula tool, swelling persisted, estimated by the EV crew at 1/32 of an inch.  When applied to gouges, bubbling of 1/16 inch was observed in the NOAX.  The material was hard to shape inside the gouge and tended to want to come out, or roll right off of the spatula.  The temperature of the test sample was recorded to be 6.7 deg C.  Testing with the sample planned for later arcjet testing on the ground had to be rescheduled due to falling behind in the EVA timeline.  The EW (Emittance Wash) application was performed on two different tile samples.  Some evidence of minor bubbling was detected.  The EW material was described as carbonated at first when coming out of the gun, however this effect quickly ceased.  The material was applied directly in the body of the tile damage crevices with the gun applicator and with a foam brush tool used to coat the edges of the tile damage.]  

Sergei and John supported the EVA with the SSRMS (Space Station Remote Manipulator System).   [First, they “walked off” of the SSRMS from MBS PDGF-1 (Mobile Service System Power & Data Grapple Fixture 1) back to the Lab PDGF.  After releasing the MBS, the now-free end of the SSRMS was maneuvered to a position where the EVA crew could tether to the SSRMS LEE (latching end effector), approximately 2 feet above the Orbiter port sill. Afterwards, the robotarm was moved to a clearance position at Orbiter aft.]

When not working with the EVA crew, Phillips and Krikalev continued cargo transfers from the MPLM.

The EVA ended at 12:36pm, with the crew repressurizing the Shuttle airlock after outer hatch closure.  Total duration: 6h 50m.  It was the 59th EVA of ISS assembly operations and the 26th from the Shuttle (33 from the ISS A/L and “Pirs” DC-1).

The CDR did the daily routine maintenance of the SM’s environment control & life support system (SOZh), including its toilet system (ASU).

Both crewmembers conducted their regular 2.5-hr. physical exercise program on the TVIS treadmill, RED resistive machine and VELO bike with bungee cord load trainer.   [Sergei’s daily protocol prescribes a strict four-day microcycle exercise with 1.5 hr on the treadmill and one hour on VELO plus load trainer (today: Day 4 of a new set).]

Afterwards, John transferred the exercise data files to the MEC (Medical Equipment Computer) for downlink, as well as the daily wristband HRM (heart rate monitor) data of the workouts on RED, followed by their erasure on the HRM storage medium (done six times a week).

ISS Location NOW

Full Size/Update
Real Time ISS TrackerMore Links

Sleep period began for all nine crewmembers again this afternoon at 3:40pm EDT.  Wakeup will be tomorrow morning at 12:10am.

Yesterday, CMG-3 experienced four instances of instabilities in a side bearing of the Hall resolver.  Since bearing instabilities can decrease the lifetime of a bearing which has inadequate lubrication, the Shuttle is controlling attitude of the mated stack (Shuttle plus ISS) until CMG-2 is brought into the steering law tomorrow.  After this time CMG-3 will remain off line for the remainder of the mated mission, and CMGs 2 & 4 (and then the new CMG 1 after EVA-2) would control attitude of the mated stack.

Also yesterday, opening the MPLM hatch proved difficult at first.  Wendy Lawrence opened the hatch MPEV (manual pressure equalization valve), but the hatch still did not move.  A few minutes later, Wendy, Sergei Krikalev and John Phillips were able to push the cover open. It is suspected that there was a higher pressure in the MPLM than in the Node, which would have made it difficult to open the hatch. 

Weekly Science Update (Expedition Eleven — 14th): 

Human Research Facility/Gas Analyzer System for Metabolic Analysis Physiology (HRF GASMAP):  Planned. 

Human Research Facility/Workstation (HRF WS):  Continuing. 

Advanced Ultrasound (ADUM):  Continuing.

Renal Stone (RS):  Continuing.

Foot/Ground Reaction Forces during Space Flight (FOOT):   The Science Officer was thanked for his fourth FOOT data collection this week. Investigators now have 8.5 hours of “great” data, and the PI considers this a very successful session. 

Space Acceleration Measurement System (SAMS):  SAMS is nominal and receiving acceleration data. 

Microgravity Acceleration Measurement System (MAMS):   MAMS remains in nominal operations. 

Protein Crystal Growth-Single Locker Thermal Enclosure System (PCG-STES): PCG-STES is performing nominally.  

Binary Colloidal Alloy Test-3 (BCAT-3):    BCAT-3 Slow Growth Sample Module will be left undisturbed in its current location by the E11 crew.  In order for the samples to potentially grow crystals that can be photographed during Increment 12 operations, the Sample Module must be left undisturbed. 

Materials ISS Experiment (MISSE):  In progress. Deployed outside on the U.S. Airlock.  Nominal and collecting data.  To be exchanged during LF-1.

Dust and Aerosol Measurement Feasibility Test (DAFT):   Nothing new.

Earth Knowledge Acquired by Middle School Students (EarthKAM):   Nothing new.

Serial Network Flow Monitor (SNFM):  Completed.

Fluid Merging Viscosity Measurement (FMVM):  Nothing new.

Space Experiment Module (SEM):  Nothing new.  Experimenters and kids are working to get the next two satchels on ULF1.1. 

Miscible Fluids in Microgravity (MFMG):   MFMG payload operations are finished. 

Educational Payload Operations (EPO):   Nothing new. 

Crew Earth Observations (CEO):   Nothing new. 

No CEO targets today.

CEO photography can be viewed and studied at the websites:

See also the website “Space Station Challenge” at:

To view the latest photos taken by the expedition 11 crew visit:

Expedition 11 Flight Crew Plans can be found at

Previous NASA ISS On-orbit Status Reports can be found here. Previous NASA Space Station Status Reports can be found here. Previous NASA Space Shuttle Processing Status Reports can be found here. A collection of all of these reports and other materials relating to Return to Flight for the Space Shuttle fleet can be found here.

ISS Altitude History

Apogee height Mean AltitudePerigee height

ISS Altitude History

For more on ISS orbit and worldwide ISS naked-eye visibility dates/times, see In addition, information on International Space Station sighting opportunities can be found at on NASA’s Human Spaceflight website. The current location of the International Space Station can be found at at NASA’s Marshall Space Flight Center. Additional satellite tracking resources can be found at

SpaceRef staff editor.